OpenStudy (anonymous):

105is what percant of 70?

OpenStudy (anonymous):

150%

OpenStudy (anonymous):

a survey at a bake offshowed that 32% of children chose strawberry cake as their favorite. A total of 750 kids took the survey. How many kids choose strawberry cake?

OpenStudy (anonymous):

32/100 is the % X/750 is the real number divide 32 by 100 then multiply by 750

OpenStudy (anonymous):

it's 240

OpenStudy (anonymous):

16 is 10% of what number?

OpenStudy (anonymous):

im i doing your homework now?! O.o it's 10/100=16/x so 10x=1600 x=160